UNDERSTANDING INTRAMOLECULAR VIBRATIONAL REDISTRIBUTION BY THE ROVIBRATIONAL ANALYSIS OF HIGH RESOLUTION INFRARED SPECTRA: THE CASE OF CHD_{2}I

\underline{CARINE MANCA TANNER}, SIEGHARD ALBERT, and MARTIN QUACK, Physical Chemistry, ETH Zurich, Switzerland.

In our group we pursue two experimental approaches to investigate Intramolecular Vibrational Redistribution (IVR)^{a}: this process can be studied by time-resolved femtosecond pump-probe experiments, or the corresponding time-dependent quantum dynamics can be obtained from stationary spectra in the IR at high frequency resolution by a time-dependent analysis using the underlying Hamiltonian and time evolution operator^{a}. Recent work in our group^{b} has shown that CH_{3}I and its deuterated isotopomers have different IVR-times, revealing different intramolecular coupling schemes for the initially excited vibrational levels. The present work is part of a larger effort to understand IVR in these molecules on the basis of high resolution spectra in the 500-12000 cm^{−1} region. In previous work we have analyzed the strong Fermi-resonance coupling between the CH-stretching and bending modes in CHD_{2}I at modest resolution, demonstrating very fast redistribution times on the order of 100 fs^{c}. We refer to this recent paper for the past literature on the topic. Here we present detailed rovibrational analysis of ν_{1} and several other fundamentals of CHD_{2}I recorded with our high resolution FTIR spectrometer Bruker ZP2001^{d} with resolutions up to 0.0008 cm^{−1}. We discuss our new results in relation to our recent work on the overtone spectra and dynamics and to the femtosecond pump-probe results.

aM. Quack, Chapter 27 in Femtosecond Chemistry, J. Manz and L. Woeste, eds. Verlag Chemie (Weinheim, 1995).

bV. Krylov, M. Nikitchenko, M. Quack, and G. Seyfang, Proc. SPIEE 5337, 178 (2004);V. Krylov, A. Kushnarenko, E. Miloglyadov, M. Quack, and G. Seyfang, Proc. SPIEE 6460, 64601D-1 (2007).

cV. Horka, M. Quack, and M. Willeke, Mol. Phys. 106, 1303 (2008).

dS. Albert and M. Quack, ChemPhysChem 8, 1271 (2007).